# Modeling Time-Variant Responses of Optical Compressors with Selective State Space Models

### v2

@@ 1 INTRODUCTION
The emulation of analog musical devices constitutes a significant portion of consumer-grade music production and performance software. The nonlinear characteristics of their electrical and electronic components give analog audio effects a distinctive timbre and sound quality that is highly sought after by both music professionals and enthusiasts. Accurate modeling of these components, individually or as a complete system, is essential for high-fidelity emulation of their sonic responses. This pursuit has become an active area of research and development, commonly known as Virtual Analog (VA) modeling.
Traditionally, analog effects are modeled using two primary approaches, physics-based and data-driven, which exhibit stark differences. The former relies on the mathematical description of physical phenomena, while the latter utilizes data measured from the actual devices. Physics-based modeling involves deriving and solving parametric mathematical equations that accurately describe an effect’s behavior and allow for manipulating the effect’s control parameters. This task is nontrivial; moreover, without simplifying assumptions, physics-based approaches can be computationally complex and may not be suitable for consumer-grade products. Conversely, data-driven approaches create digital signal processing algorithms by analyzing device measurements, aiming to replicate the input-output relationship rather than the underlying physical phenomena. These approaches often encounter limitations in flexibility. Since the model is constructed from a finite set of measurements—and considering parameter combinations are theoretically infinite—it may be confined to the work only with settings included in the measurements, or it must estimate the sound alteration process for scenarios not covered by the measurements. Generally, these models tend to exhibit much lower accuracy when predicting behaviors beyond their training data as compared to the actual device.
@@
Recent advancements in neural networks have provided novel tools to mitigate the drawbacks of existing data-driven methodology. In particular, neural networks demonstrate generalization capabilities that can surmount classical data-driven algorithms. The first successful uses of neural networks to emulate analog effects were in modeling amplifiers and distortion effects  (Mendoza,, 2005; Covert & Livingston,, 2013; Damskägg et al.,, 2019b; Wright et al.,, 2019). These works have highlighted key findings regarding the viability of artificial neural networks for Virtual Analog (VA) modeling: the models are perceptually accurate; the networks can be conditioned with the effect’s parameters, allowing for dynamic changes in the sound alteration process; accurate models can be achieved when processing raw audio samples, with the networks satisfying real-time computational constraints on consumer-grade signal processing hardware; and the networks can process single or small blocks of audio samples, thus ensuring minimal or negligible latency suitable for live audio applications.
@@
More complex types of audio effects have also been modeled using artificial neural network (ANN) modeling. This includes time-based effects such as delay  (Mikkonen et al.,, 2023) and reverb  (Ramírez et al.,, 2020), modulation effects like chorus  (Wright & Valimaki,, 2021), and dynamic effects like compressors  (Steinmetz & Reiss,, 2022; Simionato & Fasciani,, 2022, 2023, 2024). These effects typically involve a sound alteration process with complex and often lengthy temporal dependencies between the input and output signals. However, at least in their basic form, artificial neural networks are memoryless computational structures. Common variations incorporating memory, such as recurrent networks, have shown significant limitations in modeling long temporal dependencies (Bai et al.,, 2018; Simionato & Fasciani,, 2022). As a result, successful attempts to model these complex audio effects with neural networks have involved using large networks that process long blocks of audio samples. Consequently, these models generally have limited or no suitability for real-world live audio applications due to their excessive latency and high computational demands.
@@
This article introduces a method for the neural modeling of dynamic range compression (DRC), focusing on analog optical compressors. DRC is a nonlinear effect that reduces the dynamic range of the input signal  (Giannoulis et al.,, 2012). DRC is widespread in music production and live audio, typically employed to decrease the volume of louder segments, enhance quieter parts, or control transient peaks. Another common application is the ’sidechain’ mode, where an external signal triggers the compression of the input signal based on its amplitude, often used to prevent masking in concurrent sounds, such as a kick drum and bass or a lead vocal and background elements. In radio broadcasting, a compressor can automatically reduce the music volume when someone speaks.
@@
Although not all compressors feature exactly the same type and number of parameters, the function of DRC is characterized by four main controllable quantities: threshold, compression ratio, attack time, and release time. The threshold sets the amplitude level above which the compressor’s gain reduction mechanism is activated, and the ratio determines the degree of compression applied. Attack and release times dictate how quickly the gain reduction begins and ends after the threshold is exceeded. DRC applies a time-varying and level-dependent gain to the input signal. The behavior of DRCs is described by their compression curves, which map input amplitudes to output amplitudes. Below the threshold level, the curve represents a one-to-one relationship between input and output amplitudes, followed by a gain reduction part above the threshold. The compression curve features a ’knee’ at the threshold amplitude; the smoothness of this knee — which determines how gradually the compression is applied — varies between devices in analog units due to their circuitry and is adjustable in most digital compressors. However, the compression curve is dynamic rather than static due to the non-zero attack and release times. As a result, the actual sound alteration process smoothly transitions from a linear relationship to the nonlinear compression curve, influenced by the duration set by the attack parameter, and then gradually returns to the original dynamics dictated by the release parameter.
The variable attack and release times help to adapt the compression to the input signal and the specific aesthetic needs. Faster attack times result in quick compression, which is useful for the immediate taming of peaks, while longer attack times allow for smoother compression and the preservation of some peak characteristics. Additionally, the compressor’s design determines the minimum and maximum allowable times for compression. For example, field-effect transistor (FET) based compressors can effectively handle particularly sharp amplitude envelopes due to their quick attack and release times.
@@
Recurrent Neural Networks (RNNs) and State Space Models (SSMs) have demonstrated their capability to track the history of inputs across various application fields. However, when processing raw audio samples at a sampling rate of $48$ kHz, managing a time dependency as brief as $100$ ms requires the network to remember events that occurred $4,800$ iterations earlier, assuming the network predicts one audio sample per inference iteration. This requirement presents the challenge of incorporating an internal mechanism capable of maintaining relevant input information in the model’s memory for a considerable number of iterations. Addressing variable long-term dependencies, such as those encountered in the case of audio compression, is a significant challenge.
Building upon our previous contributions to analog DRC modeling, we have enhanced the recurrent neural architecture we previously proposed by incorporating the Selective State Space (S6) model (Gu & Dao,, 2023) while preserving the original design’s low latency and small computational costs. State space models have demonstrated their potential in sequence-to-sequence modeling tasks and, more recently, in capturing the significant temporal characteristics of analog effects  (Simionato & Fasciani,, 2024). Specifically, the Structured State Space (S4D) model  (Gu et al.,, 2022b) has successfully applied to optical DRC.
However, a disadvantage of the S4D model is that the parameters governing the computation of the internal states are not influenced by the input, which can be a limitation when the input plays a crucial role in the device’s operation. In optical DRC, the input signal significantly affects the applied compression. The more recent S6 model overcomes this issue, and we demonstrate how this adaptation improves the learning process for analog optical compressor devices. We further enhanced the black-box modeling approach by integrating all the user-controllable parameters using Feature-wise Linear Modulation (FiLM)  (Perez et al.,, 2018) and Gated Linear Units (GLU)  (Dauphin et al.,, 2017) methods to condition the networks. Following the proposal in  (Fasciani et al.,, 2024a), we placed the conditioning layer after the State Space layer. We also introduced a second layer after the conditioning block to provide additional handling of nonlinearities and time dependencies. Given that the DRC’s control parameters govern various aspects of its operation, we divided the conditioning vector into two parts: one addressing dynamics and another focusing on timing characteristics, for which we employed the Temporal FiLM method  (Birnbaum et al.,, 2019). Furthermore, we confirmed the efficacy of using the window technique, which aids the network in inferring the correct output sample by providing auxiliary past input information. In this context, we leveraged the spectral magnitude information from the input signal to facilitate the conditioning process.
We compare the proposed modeling method with other approaches based on the same architecture but featuring different types of layers: Long Short-Term Memory (LSTM)  (Hochreiter & Schmidhuber,, 1997) and S4D. We evaluate the models’ overall learning abilities, focusing on extreme scenarios such as maximum compression and the slowest attack and release times. Additionally, we assess the ability of the models to predict compression for unseen combinations of conditioning parameters. We investigate the impact of the control parameters’ dataset sampling resolution on the network’s ability to interpolate between values seen during training. Finally, listening tests are conducted to evaluate the perceptual quality of the models subjectively.
@@ 2 BACKGROUND
Steinmetz et al., in what appears to be the earliest attempt to model a DRC  (Steinmetz & Reiss,, 2022), proposed the use of Temporal Convolutional Networks (TCNs). In this work, the modeled device is the Teletronix LA-2A leveling amplifier, which features an optical gain reduction mechanism - challenges associated with its modeling were discussed in the introductory section. The model operates at a sampling rate of $44.1$ kHz. The TCN employs dilated convolutional layers with rapidly growing dilation factors ranging from $2^{2}$ to $02^{1}$ 0. As the dilation factor increases, so does the receptive field, enabling the network to handle long temporal dependencies without adding more layers. This approach led to the development of a model with relatively low computational demands; however, it still required significantly long segments of the input signal, which is unfavorable for latency. The various models used in comparative experiments exhibit input-output latencies between $100$ and $1000$ ms, corresponding to their respective receptive field lengths. TCN block used to model the LA-2A consists of a convolutional layer, batch normalization, conditioning using FiLM, and the Parametric Rectified Linear Unit (PReLU) activation function. The authors compared networks with $4$, $5$, $10$ TCN blocks, with each convolutional layer having $32$ channelsand with kernel sizes of $5$, $13$, and $15$, respectively. The models correspondingly produced receptive fields of $100$ ms, $300$ ms, and $1000$ ms. The model with a receptive field of $300$ ms demonstrated the best accuracy.
@@
To improve the networks’ ability to encode the past signal information without relying on large layers, we proposed an Encoder-Decoder (ED) architecture initially utilized to model the optical compressor TubeTech CL 1B  (Simionato & Fasciani,, 2022). The network is structured with two LSTM layers; the first functions as an encoder, processing a short segment of past input samples alongside the compressor control parameters, while the second acts as a decoder. The decoder uses either the current input sample for single-sample prediction at each inference iteration or a current block of input samples for block-wise prediction. As the encoder processes the input signal, it computes and updates its internal states, which are then utilized as initial states by the LSTM in the decoder. This allows the states to effectively encapsulate a summary of the encoder’s input signal, aiding the decoder in the output inference. This architecture demonstrated a significant improvement in accuracy compared to LSTMs and FC networks with an identical number of trainable parameters. The model operates at a sampling rate of $48$ kHz and incorporates only two variable control parameters, namely threshold and ratio. This study investigated networks with a maximum of $128$ units per layer and short input segments, ranging from $2$ to $16$ audio samples, resulting in a worst-case latency of $0.33$ ms. The results show a positive correlation between segment size and accuracy.
@@
Subsequently, we augmented the model by including attack and release times as variable parameters with an enhanced architecture, where a convolutional layer replaced the encoder’s LSTM  (Simionato & Fasciani,, 2023). This modification led to improved accuracy, albeit with a minor increase in the computational cost. However, this approach eliminated the continuous updating of internal states across sequences; states are recalculated at each iteration based on the number of the encoder’s input samples. We experimented with up to $64$ samples, bounding the worst-case latency to $1.33$ ms. We evaluated $32$, $64$, and $128$ as model input segment sizes, with the smallest delivering the best accuracy. Network size was kept small to limit computational complexity, with a maximum of $128$ units per layer, consistent with prior work. The best-performing model from our approach was compared against the top TCN model from  (Steinmetz & Reiss,, 2022) using our TubeTech CL 1B dataset and their Teletronix LA-2A dataset. Results showed that our model exhibited better accuracy in the CL 1B case. In comparison, the accuracy was slightly worse but comparable to that of the TCN in the LA-2A case despite the TCN model using significantly larger input segments. However, when the TCN was adapted to work with shorter input segments—thereby matching the input-output latency of our model—its accuracy was significantly inferior to that of our ED model. Moreover, the CL1B dataset included cases with significant compression at low thresholds (e.g., $-30$, $-40$ dBu), which results in audible artifacts due to minor signal amplitude mismatching at the edges of consecutive output audio segments. These artifacts were observed in both the ED and TCN models. However, the models exhibit limitations in scenarios of heavy compression, where they fail to replicate the abrupt and rapid gain reduction that typically occurs with onsets. Rapid attacks, when combined with heavy compression, are particularly challenging to model, as are extended-release times, which can span up to $10$ seconds in the case of the CL 1B.
@@
A gray-box model, which combines information about the device’s functioning with automatic parameter estimation, was proposed for the Teletronix LA-2A optical leveling amplifier  (Wright & Välimäki,, 2022). The model incorporates the structure of a traditional digital compressor and predicts the parameters of the compression curve, including the threshold, ratio, and knee, using FC- and RNN-based networks. Similar techniques are employed to predict the timing parameters that define the attack and release envelopes of the equations designed and used to describe the compressor. This approach leverages domain knowledge to help generate audio with fewer artifacts and reduce the number of trainable parameters. However, the accuracy of this approach is constrained by the inherent approximations present in the equations representing the system. In this study, for example, timing effects are modeled using three methods: the first assumes equal attack and release times, independent of the input signal; the second assumes varying attack and release times, but still not in response to the input signal; and the third assumes attack and release times are equal but do vary based on the input signal. The latter two methods demonstrated better performance. However, all these scenarios fail to accurately represent the actual timing behavior of the device, as, in reality, attack and release times may vary independently of each other and are influenced by the signal’s energy. Consequently, due to these approximations, the model is inaccurately biased, which compromises its precision. On the other hand, the model brings the advantage of requiring $10\%$ of the number of operations needed by a black-box GRU model with $32$ units. In addition, the gray-box model does not require long input segments to achieve greater accuracy and, as a result, does not introduce added latency.
@@
In most optical compressors, such as the LA-2A and CL1B, formalizing the variable attack and release behavior with mathematical expressions is not straightforward, which makes the use of black-box design potentially advantageous. The complex nonlinear and time-variant behavior of compressors presents a significant challenge for traditional RNN and CNN architectures. Recent advances in sequence-to-sequence modeling have introduced State Space models (SSMs) that rely on a continuous-time representation of the state space formulation  (Gu et al.,, 2022a). These models have demonstrated promising results, especially with long sequences. Similar to RNNs, they retain state information while addressing the vanishing memory problem by eliminating nonlinearity in the computation of states. We implemented a variant of SSMs, the S4D model, for a range of analog effects, including the aforementioned compressors  (Simionato & Fasciani,, 2024). The results indicated that the S4D model improved the modeling of compressor effects but did not markedly outperform LSTMs. A limitation of these models became apparent when the parameters were found to be static and not influenced by the input signal  (Gu & Dao,, 2023). Incorporating input-dependent variation can enhance the model’s content-based predictions, a critical aspect for compressors where the degree of compression is dependent on the present and past input signal’s level.
It is also worth mentioning that prior to the advent of SSM, transformer architectures had already shown improvements in multiple application domains over CNNs and RNNs. However, the context of the input window also affects the performance of attention operations in transformers. Therefore, just like CNNs, these architectures require as many past input samples as necessary to account for temporal dependencies in the input-output data. Due to this characteristic, which is detrimental to latency, they have not been employed in the modeling of analog effects.

### v3

@@ 1 INTRODUCTION
The emulation of analog musical devices constitutes a significant portion of consumer-grade music production and performance software. The nonlinear characteristics of their electrical and electronic components give analog audio effects a distinctive timbre and sound quality that is highly sought after by both music professionals and enthusiasts. Accurate modeling of these components has become an active area of research and development, commonly known as Virtual Analog (VA) modeling.
Modeling of analog effects typically employs one of two main approaches: physics-based or data-driven, each with distinct characteristics. The physics-based method depends on mathematical descriptions of physical processes, whereas the data-driven strategy leverages data collected from the devices themselves. Physics-based modeling involves deriving and solving parametric mathematical equations that accurately describe an effect’s behavior and allow for manipulating the effect’s control parameters. This task is nontrivial; moreover, without simplifying assumptions, physics-based approaches can be computationally complex and may not be suitable for consumer-grade products. Among physics-based approaches, wave digital filter (WDF) techniques have gained particular popularity; these have been utilized for modeling vacuum tube-based circuits  (Dunkel et al.,, 2016), envelope filter guitar effect circuit consisting of operational transconductance amplifiers  (Bogason & Werner,, 2017), cascaded amplifiers  (Zhang & Smith,, 2018), a booster guitar pedal based on field-effect transistors  (Bogason,, 2018), ring modulators  (Bernardini et al.,, 2018), and the phaser MXR 90  (Giampiccolo et al.,, 2024).
Conversely, data-driven approaches derive digital signal processing algorithms by analyzing device measurements, aiming to replicate the input-output relationship rather than the underlying physical phenomena. These approaches often encounter limitations in terms of flexibility. Since models are obtained from a limited set of measurements, and there is a theoretically infinite combination of possible inputs, they may either be limited to operating exclusively within scenarios represented in the measurements, where sufficient accuracy can be verified, or they can be used to predict outputs for input combinations unseen in the measurement data, where accuracy may not be absolutely guaranteed, especially when inferring values outside the range of the measurement data. Common examples of data-driven approaches include block-oriented models such as the Hammerstein and Wiener models, as well as their various combinations  (Schoukens & Tiels,, 2017). These models have been utilized for guitar distortion pedals  (Eichas et al.,, 2015), guitar amplifiers  (Schuck Jr & Bodmann,, 2016; Eichas et al.,, 2017), distortion circuits  (Eichas & Zölzer,, 2016), and guitar overdrive effect pedals  (Köper & Holters,, 2020).
@@
Recent advancements in neural networks have provided novel tools to mitigate the drawbacks of existing data-driven methodology. In particular, neural networks demonstrate generalization capabilities that can surmount classical data-driven algorithms. The first successful uses of artificial neural networks (ANNs) to emulate analog effects were in modeling amplifiers and distortion effects  (Mendoza,, 2005; Covert & Livingston,, 2013; Damskägg et al.,, 2019b; Wright et al.,, 2019). These works have highlighted key findings regarding the viability of artificial neural networks for Virtual Analog (VA) modeling: the models are perceptually accurate; the networks can be conditioned with the effect’s parameters, allowing for dynamic changes in the sound alteration process; accurate models can be achieved when processing raw audio samples, with the networks satisfying real-time computational constraints on consumer-grade signal processing hardware; and the networks can process single or small blocks of audio samples, thus ensuring minimal or negligible latency suitable for live audio applications.
@@
More complex types of audio effects have also been successfully modeled using ANN-based modeling techniques. This includes time-based effects such as delay  (Mikkonen et al.,, 2023) and reverb  (Ramírez et al.,, 2020), modulation effects like chorus  (Wright & Valimaki,, 2021), and dynamic effects like compressors  (Steinmetz & Reiss,, 2022; Simionato & Fasciani,, 2022, 2023). These effects typically involve a sound alteration process with complex and often lengthy temporal dependencies between the input and output signals. However, at least in their basic form, artificial neural networks are memoryless computational structures. Common variations incorporating memory, such as recurrent networks, have shown significant limitations in modeling long temporal dependencies  (Bai et al.,, 2018; Simionato & Fasciani,, 2022). As a result, successful attempts to model these complex audio effects with ANNs have utilized large-scale networks that process audio signals in large blocks of samples. Consequently, these models generally have poor suitability for actual live audio applications due to their excessive latency or high computational complexity.
@@
This article introduces a method for the neural modeling of dynamic range compression (DRC), focusing on analog optical compressors. DRC is a nonlinear effect that reduces the dynamic range of the input signal  (Giannoulis et al.,, 2012). DRC is widespread in music production and live audio, typically employed to decrease the volume of louder segments, enhance quieter parts, or control transient peaks. Another common application is the ‘sidechain’ mode, where an external signal triggers the compression of the input signal based on its amplitude, often used to prevent masking in concurrent sounds, such as a kick drum and bass or a lead vocal and background elements. In radio broadcasting, a compressor can automatically reduce the music volume when someone speaks.
@@
Although not all compressors feature the same type and number of parameters, the function of DRC is characterized by four main controllable quantities: threshold, compression ratio, attack time, and release time. The threshold sets the amplitude level above which the compressor’s gain reduction mechanism is activated, and the ratio determines the degree of compression applied. Attack and release times dictate how quickly the gain reduction begins and ends after exceeding the threshold. DRC applies a time-varying and level-dependent gain to the input signal.
@@
Building upon our previous contributions to analog DRC modeling, we have enhanced the neural architecture incorporating the Selective State Space (S6) model (Gu & Dao,, 2023) while preserving the original design’s low latency and small computational costs. State space models have demonstrated their potential in sequence-to-sequence modeling tasks and, more recently, in capturing the significant temporal characteristics of analog effects  (Simionato & Fasciani,, 2024).
We further enhanced the black-box modeling approach by integrating all the user-controllable parameters using Feature-wise Linear Modulation (FiLM)  (Perez et al.,, 2018) and Gated Linear Units (GLU)  (Dauphin et al.,, 2017) methods to condition the networks. Following the proposal in  (Fasciani et al.,, 2024a), we placed the conditioning layer after the State Space layer. We also introduced a second layer after the conditioning block to provide additional handling of nonlinearities and time dependencies. Given that the DRC’s control parameters govern various aspects of its operation, we divided the conditioning vector into two parts: one addressing dynamics and another focusing on timing characteristics, for which we employed the Temporal FiLM method  (Birnbaum et al.,, 2019). Furthermore, we confirmed the efficacy of using the window technique, which aids the network in inferring the correct output sample by providing auxiliary past input information. In this context, we leveraged the spectral magnitude information from the input signal to facilitate the conditioning process.
We compare the proposed modeling method with other approaches based on the same architecture but featuring different types of layers: Long Short-Term Memory (LSTM)  (Hochreiter & Schmidhuber,, 1997) and Structured State Space (S4D) model  (Gu et al.,, 2022b). We evaluate the models’ overall learning abilities and the ability of the models to predict compression for unseen combinations of conditioning parameters. We investigate the impact of the control parameters’ dataset sampling resolution on the network’s ability to interpolate between values seen during training. Finally, listening tests are conducted to evaluate the perceptual quality of the models subjectively.
@@ 2 BACKGROUND
Steinmetz et al., in what appears to be the earliest attempt to model a DRC  (Steinmetz & Reiss,, 2022), proposed the use of Temporal Convolutional Networks (TCNs). In this work, the modeled device is the Teletronix LA-2A leveling amplifier, which features an optical gain reduction mechanism - challenges associated with its modeling were discussed in the introductory section. The model operates at a sampling rate of $44.1$ kHz. The TCN employs dilated convolutional layers with rapidly growing dilation factors ranging from $2^{2}$ to $2^{10}$. As the dilation factor increases, so does the receptive field, enabling the network to handle long temporal dependencies without adding more layers. This approach led to the development of a model with relatively low computational demands; however, it still required significantly long segments of the input signal, which is unfavorable for latency. The various models used in comparative experiments exhibit input-output latencies between $100$ and $1000$ ms, corresponding to their respective receptive field lengths. The model with a receptive field of $300$ ms demonstrated the best accuracy.
@@
To enhance past signal information encoding in networks without increasing layer size, we proposed an Encoder-Decoder (ED) architecture utilized to model the optical compressor TubeTech CL 1B  (Simionato & Fasciani,, 2022). The network comprises two LSTM layers: an encoder working with a brief history of input samples and compressor settings, and a decoder for outputting predictions per input sample. The encoder continuously computes its internal states, which are then utilized by the decoder’s LSTM layer as initial states. This transfer facilitates the decoder in generating accurate predictions by making use of a condensed history of past inputs. This ED approach notably outperformed traditional LSTMs and FC networks in accuracy, despite having a similar count of trainable parameters. Operating at a sampling rate of $48$ kHz, the model incorporates only two variable control parameters: threshold and ratio. This study examined networks with input segments ranging from $2$ to $16$ audio samples, resulting in a worst-case latency of $0.33$ ms. The results show a positive correlation between input segment size and accuracy.
@@
Subsequently, we augmented the model by including attack and release times as variable parameters with an enhanced architecture, where a convolutional layer replaced the encoder’s LSTM  (Simionato & Fasciani,, 2023). This modification led to improved accuracy, albeit with a minor increase in the computational cost. However, this approach eliminated the continuous updating of internal states across sequences; states are recalculated at each iteration based on the number of the encoder’s input samples. We experimented with up to $64$ samples, bounding the worst-case latency to $1.33$ ms. The network size was kept small to limit computational complexity.
The best-performing model among those developed was compared against the top TCN model from  (Steinmetz & Reiss,, 2022) using our TubeTech CL 1B dataset and their Teletronix LA-2A dataset. Results showed that our model exhibited better accuracy in the CL 1B case. In comparison, the accuracy was slightly worse but comparable to that of the TCN in the LA-2A case despite the TCN model using significantly larger input segments. However, when the TCN was adapted to work with shorter input segments—thereby matching the input-output latency of our model—its accuracy was significantly inferior to that of our ED model. Moreover, the CL1B dataset included cases with significant compression at low thresholds (e.g., $-30$, $-40$ dBu), which results in audible artifacts due to minor signal amplitude mismatching at the edges of consecutive output audio segments. These artifacts were observed in both the ED and TCN models. However, the models exhibit limitations in scenarios of heavy compression, where they fail to replicate the abrupt and rapid gain reduction that typically occurs with onsets. Rapid attacks, when combined with heavy compression, are particularly challenging to model, as are extended-release times, which can span up to $10$ seconds in the case of the CL 1B.
@@
A gray-box model, which combines information about the device’s functioning with automatic parameter estimation, was proposed for the Teletronix LA-2A optical leveling amplifier  (Wright & Välimäki,, 2022). The model incorporates the structure of a traditional digital compressor and predicts the parameters of the compression curve, including the threshold, ratio, and knee, using FC- and RNN-based networks. Similar techniques are employed to predict the timing parameters that define the attack and release envelopes of the equations designed and used to describe the compressor. This approach leverages domain knowledge to help generate audio with fewer artifacts and reduce the number of trainable parameters.
The accuracy of this approach, however, is constrained by the inherent approximations present in the equations representing the system. In this study, for example, timing effects are modeled using three methods: the first assumes equal attack and release times, independent of the input signal; the second assumes varying attack and release times, but still not in response to the input signal; and the third assumes attack and release times are equal but do vary based on the input signal. The latter two methods demonstrated better performance. However, all these scenarios fail to accurately represent the actual timing behavior of the device, as, in reality, attack and release times may vary independently of each other and are influenced by the signal’s energy. Consequently, due to these approximations, the model is inaccurately biased, which compromises its precision. On the other hand, the model brings the advantage of requiring $10\%$ of the number of operations needed by a black-box GRU model with $32$ units. In addition, the gray-box model does not require long input segments to achieve greater accuracy and, as a result, does not introduce added latency.
@@
In most optical compressors, such as the LA-2A and CL1B, formalizing the variable attack and release behavior with mathematical expressions is not straightforward, which makes the use of black-box design potentially advantageous. The complex nonlinear and time-variant behavior of compressors presents a significant challenge for traditional RNN and CNN architectures. Recent advances in sequence-to-sequence modeling have introduced State Space models (SSMs) that rely on a continuous-time representation of the state space formulation  (Gu et al.,, 2022a). These models have demonstrated promising results, especially with long sequences. Similar to RNNs, they retain state information while addressing the vanishing memory problem by eliminating nonlinearity in the computation of states.